# Chain-of-Influence: Tracing Interdependencies Across Time and Features in Clinical Predictive Modelings

### v1

@@ 6 Limitations
7 Conclusion
In summary, the Chain-of-Influence framework advances the state-of-the-art in clinical predictive modeling by bridging the gap between performance and interpretability. Our approach explicitly models and visualizes the temporal and cross-feature dependencies that underlie complex clinical outcomes, achieving robust performance gains across heterogeneous datasets while providing clinicians with actionable, transparent explanations. CoI’s dual attention and influence tracing mechanism transforms black-box deep learning predictions into interpretable, dynamic influence maps that mirror clinical reasoning and inform optimal intervention strategies.
Looking forward, the flexibility of the Chain-of-Influence paradigm suggests broad applicability across diverse temporal and multi-modal clinical scenarios. Future work will explore extending the framework to handle richer data modalities, finer-grained temporal granularity, and integration with real-time clinical decision systems. By providing a transparent and generalizable approach to understanding the interplay of clinical variables over time, CoI represents a critical step toward safer, more accountable, and more effective deployment of AI in healthcare.
@@ A.2 MIMIC-IV Dataset Characteristics
Appendix B Data Processing Implementation Details
B.1 MIMIC-IV Preprocessing Pipeline
The preprocessing pipeline for MIMIC-IV data involves several computational steps designed to handle the scale and complexity of the raw database:
Computational Requirements:
  • Memory usage: ~32GB RAM for full dataset processing
  • Processing time: ~4-6 hours on 16-core CPU
  • Storage: Raw data (120GB), processed data (8.2GB)
Batch Processing Strategy: To manage memory constraints, patient processing is batched in groups of 1,000 patients, with relevant chartevents and labevents filtered for each batch to minimize memory footprint.
Temporal Alignment: All timestamps are aligned to ICU admission time (intime), with measurements indexed by hour offset (0-47) from admission. This standardization enables consistent temporal modeling across patients.
B.2 Feature Engineering Considerations
Clinical Rationale for Feature Selection:
  • Vital Signs: Selected based on immediate availability and prognostic value in ICU mortality prediction
  • Laboratory Tests: Chosen to represent major organ systems (kidney, metabolism, hematology, infection)
  • Exclusions: Medication and intervention data excluded to focus on physiological state rather than treatment effects
Temporal Binning Strategy: Measurements within each hour are aggregated using median values to reduce noise while preserving clinical trends. This approach balances temporal resolution with data stability.
Appendix C Hyperparameter Configuration Details
This appendix provides comprehensive hyperparameter details for four key models evaluated on the Chronic Kidney Disease (CKD) progression prediction task: BiLSTM, RETAIN, StageNet, and Chain-of-Influence (CoI). For each model, we present both the search space of hyperparameter candidates and the optimal configuration identified through hyperparameter optimization.
C.1 BiLSTM (Bidirectional Long Short-Term Memory)
@@
Table 7: BiLSTM hyperparameter search space (Total: $3\times 3\times 3\times 3\times 3=243$ combinations)
@@
Table 8: BiLSTM optimal hyperparameters
C.2 RETAIN (REverse Time AttentIoN)
@@
Batch Size {16, 32, 64}
Dropout Rate {0.1, 0.2, 0.3}
Embedding Dimension {32, 64, 128}
Hidden Dimension {64, 128, 256}
@@
Number of Attention Heads {4, 8, 16}
@@
Table 9: RETAIN hyperparameter search space (Total: $3\times 3\times 3\times 3\times 3\times 3\times 3=2187$ combinations)
@@
Table 10: RETAIN optimal hyperparameters
C.3 StageNet (Stage-aware Neural Network)
@@
Table 11: StageNet hyperparameter search space (Total: $3\times 3\times 3\times 3\times 3\times 3=729$ combinations)
@@
Table 12: StageNet optimal hyperparameters
C.4 CoI (Chain-of-Influence with Dynamic Tanh Normalization)
@@
Number of Layers {2, 4, 6}
Alpha Initialization {0.8} (DyT-specific)
Table 13: CoI hyperparameter search space (Total: $1\times 3\times 3\times 3\times 2\times 3\times 3\times 1=486$ combinations)
@@
Table 14: CoI optimal hyperparameters
C.5 Training Configuration
All models were trained with the following common settings:
@@ C.5 Training Configuration
Early Stopping Patience 7
@@
Validation Split 20%
Test Split 20%
Random Seed 42
Table 15: Common training configuration for all models

### v2

@@ A.2 MIMIC-IV Dataset Characteristics
Appendix B Hyperparameter Configuration Details
This appendix provides comprehensive hyperparameter details for five key models evaluated on the Chronic Kidney Disease (CKD) progression prediction task: BiLSTM, RETAIN, StageNet, BEHRT-style Transformer, and Chain-of-Influence (CoI). For each model, we present both the search space of hyperparameter candidates and the optimal configuration identified through hyperparameter optimization.
B.1 BiLSTM (Bidirectional Long Short-Term Memory)
@@
Table 9: BiLSTM hyperparameter search space (Total: $3\times 3\times 3\times 3\times 3=243$ combinations).
@@
Table 10: BiLSTM optimal hyperparameters.
B.2 RETAIN (REverse Time AttentIoN)
@@
Batch Size {32, 64, 128}
Dropout Rate {0.0, 0.1, 0.2}
Embedding Dimension {16, 32, 64}
Hidden Dimension {32, 64, 128}
@@
Number of Attention Heads {2, 4, 8, 16}
@@
Table 11: RETAIN hyperparameter search space (Total: $3^{6}\times 4=2916$ combinations).
@@
Table 12: RETAIN optimal hyperparameters.
B.3 StageNet (Stage-aware Neural Network)
@@
Table 13: StageNet hyperparameter search space (Total: $3^{6}=729$ combinations).
@@
Table 14: StageNet optimal hyperparameters.
B.4 BEHRT-style Transformer
@@
Batch Size {32, 64, 128}
Dropout Rate {0.1, 0.2, 0.3}
Embedding Dimension {16, 32, 64}
Hidden Dimension {64, 128, 256}
Learning Rate {0.0001, 0.001}
Number of Attention Heads {2, 4, 8}
Number of Layers {2, 4, 6}
Table 15: BEHRT-style Transformer hyperparameter search space (Total: $3\times 3\times 3\times 3\times 2\times 3\times 3=1458$ combinations).
Optimal Configuration
Hyperparameter Best Value
@@
Dropout Rate 0.2
Embedding Dimension 32
Hidden Dimension 128
Learning Rate 0.0001
Number of Attention Heads 4
Number of Layers 4
Table 16: BEHRT-style Transformer optimal hyperparameters.
B.5 CoI (Chain-of-Influence with Dynamic Tanh Normalization)
Hyperparameter Search Space
Hyperparameter Candidates
Batch Size {32,64,128}
@@ B.5 CoI (Chain-of-Influence with Dynamic Tanh Normalization)
Number of Layers {4, 6, 8}
Alpha Initialization (DyT) {0.2, 0.4, 0.6, 0.8}
Table 17: CoI hyperparameter search space (Total: $3\times 3\times 3\times 3\times 2\times 3\times 3\times 4=5832$ combinations).
@@
Table 18: CoI optimal hyperparameters.
B.6 Training Configuration
All models were trained under a common protocol:
@@ B.6 Training Configuration
Early Stopping Patience 10 (on validation F1 )
@@
Validation Split 20% of training data
Test Split 20% (held-out)
Number of Random Seeds 5
Table 19: Common training configuration for all models.
Appendix C Training Details
C.1 Computational Cost and Training Time
[…]
Model CKD (hours/run) MIMIC-IV (hours/run)
BiLSTM 0.10 0.25
RETAIN 0.15 0.40
StageNet 0.30 0.90
BEHRT-style Transformer 0.35 1.00
CoI 0.40 1.30
[…]
C.2 Comparison of Class-Imbalance Strategies
Both datasets exhibit substantial class imbalance—CKD progression (6.0% positive class) and MIMIC-IV mortality (10.8% positive class)—which can significantly affect both optimization and calibration of predictive models. To justify our imbalance-handling approach, we compare several commonly used strategies applied to CoI on both datasets:
  • None: no explicit rebalancing beyond the original class distribution.
  • SMOTE: standard Synthetic Minority Oversampling Technique applied at the patient-level.
  • ADASYN: adaptive synthetic oversampling that focuses on harder minority examples.
  • ENN: Edited Nearest Neighbors undersampling of majority examples near the decision boundary.
  • SMOTE+ENN: sequential SMOTE oversampling followed by ENN undersampling.
  • Weighted Loss: class-weighted BCEWithLogitsLoss with inverse-frequency weights.
  • TSMOTE: temporal SMOTE variant that interpolates entire minority trajectories while preserving time ordering.
Table 21 reports AUROC and F1 for each strategy on CKD. All results are averaged over five random seeds using the same architecture and optimization settings as in the main CoI experiments.
Imbalance Strategy CKD Dataset MIMIC-IV Dataset
[…]
None 0.930 $\pm$ 0.020 0.580 $\pm$ 0.030 0.920 $\pm$ 0.018 0.780 $\pm$ 0.025
SMOTE 0.945 $\pm$ 0.018 0.640 $\pm$ 0.026 0.935 $\pm$ 0.016 0.820 $\pm$ 0.022
ADASYN 0.942 $\pm$ 0.019 0.630 $\pm$ 0.028 0.932 $\pm$ 0.017 0.815 $\pm$ 0.023
ENN 0.938 $\pm$ 0.022 0.600 $\pm$ 0.027 0.928 $\pm$ 0.019 0.800 $\pm$ 0.024
SMOTE+ENN 0.952 $\pm$ 0.016 0.690 $\pm$ 0.022 0.942 $\pm$ 0.014 0.840 $\pm$ 0.020
Weighted Loss 0.955 $\pm$ 0.015 0.700 $\pm$ 0.020 0.945 $\pm$ 0.013 0.850 $\pm$ 0.018
TSMOTE (ours) 0.960 $\pm$ 0.010 0.721 $\pm$ 0.013 0.950 $\pm$ 0.009 0.865 $\pm$ 0.012
Table 21: Comparison of class-imbalance handling strategies for CoI on both datasets. Values denote mean $\pm$ 95% confidence intervals over 5 runs. Naive training without rebalancing (None) performs substantially worse with higher variance, while TSMOTE yields the best F1-AUROC trade-off on both datasets and is therefore used in all main experiments.
Appendix D Experimental Result Details
D.1 Performance comparison details across datasets
CKD Dataset
[…]
Bi-LSTM 0.910 $\pm$ 0.012 0.616 $\pm$ 0.018 0.900 $\pm$ 0.011 0.700 $\pm$ 0.016 0.550 $\pm$ 0.018
RETAIN 0.930 $\pm$ 0.011 0.671 $\pm$ 0.017 0.920 $\pm$ 0.010 0.760 $\pm$ 0.015 0.600 $\pm$ 0.017
StageNet 0.940 $\pm$ 0.010 0.685 $\pm$ 0.016 0.930 $\pm$ 0.010 0.780 $\pm$ 0.014 0.610 $\pm$ 0.016
BEHRT-style 0.955 $\pm$ 0.010 0.715 $\pm$ 0.015 0.954 $\pm$ 0.009 0.810 $\pm$ 0.013 0.640 $\pm$ 0.017
CoI 0.960 $\pm$ 0.009 0.721 $\pm$ 0.013 0.950 $\pm$ 0.008 0.820 $\pm$ 0.011 0.660 $\pm$ 0.016
MIMIC-IV Dataset
[…]
Bi-LSTM 0.700 $\pm$ 0.018 0.267 $\pm$ 0.018 0.890 $\pm$ 0.012 0.400 $\pm$ 0.017 0.200 $\pm$ 0.018
RETAIN 0.944 $\pm$ 0.013 0.667 $\pm$ 0.018 0.900 $\pm$ 0.011 0.500 $\pm$ 0.016 1.000 $\pm$ 0.000
StageNet 0.920 $\pm$ 0.012 0.768 $\pm$ 0.017 0.930 $\pm$ 0.010 0.700 $\pm$ 0.015 0.850 $\pm$ 0.017
BEHRT-style 0.948 $\pm$ 0.011 0.834 $\pm$ 0.016 0.945 $\pm$ 0.009 0.790 $\pm$ 0.014 0.884 $\pm$ 0.017
CoI 0.950 $\pm$ 0.010 0.865 $\pm$ 0.015 0.950 $\pm$ 0.009 0.850 $\pm$ 0.013 0.880 $\pm$ 0.016
Table 22: Performance comparison on CKD and MIMIC-IV datasets. Values denote mean $\pm$ 95% confidence intervals over 5 runs (per-metric standard deviations <0.015<0.015). Best results per column are highlighted in bold.
D.2 Ablation Study Result Details
Model Variant CKD Dataset MIMIC-IV Dataset
[…]
CoI (Full) 0.960 $\pm$ 0.010 - 0.721 $\pm$ 0.013 - 0.950 $\pm$ 0.009 - 0.865 $\pm$ 0.012 -
w/o Temporal Attn 0.910 $\pm$ 0.011 -0.050 0.650 $\pm$ 0.014 -0.071 0.860 $\pm$ 0.012 -0.090 0.750 $\pm$ 0.015 -0.115
w/o Feature Attn 0.895 $\pm$ 0.010 -0.065 0.630 $\pm$ 0.015 -0.091 0.895 $\pm$ 0.011 -0.055 0.810 $\pm$ 0.013 -0.055
w/o Transformer 0.885 $\pm$ 0.012 -0.075 0.620 $\pm$ 0.015 -0.101 0.870 $\pm$ 0.013 -0.080 0.765 $\pm$ 0.016 -0.100
Temporal only 0.900 $\pm$ 0.011 -0.060 0.640 $\pm$ 0.013 -0.081 0.905 $\pm$ 0.010 -0.045 0.820 $\pm$ 0.012 -0.045
Feature only 0.915 $\pm$ 0.009 -0.045 0.660 $\pm$ 0.014 -0.061 0.865 $\pm$ 0.012 -0.085 0.760 $\pm$ 0.015 -0.105
w/ LayerNorm 0.940 $\pm$ 0.010 -0.020 0.695 $\pm$ 0.012 -0.026 0.925 $\pm$ 0.011 -0.025 0.835 $\pm$ 0.013 -0.030
Transformer Only 0.875 $\pm$ 0.013 -0.085 0.605 $\pm$ 0.016 -0.116 0.845 $\pm$ 0.014 -0.105 0.730 $\pm$ 0.017 -0.135
Table 23: Ablation study results on CKD and MIMIC-IV datasets. Values denote mean $\pm$ 95% confidence intervals over 5 runs. $\Delta$ shows absolute performance drop from the full CoI model.
Appendix E Chain-of-Influence Visualizations